# The Loss of Control Playbook: Degrees, Dynamics, and Preparedness

### v1

@@ 2.1 Literature Reviewed
Hastings-Woodhouse, Sarah. 2024. ”Could we switch off a dangerous AI?” https://futureoflife.org/ai/could-we-switch-off-a-dangerous-ai/.
@@
Schmitz, Chris, Jonathan Rystrøm, Jan Batzner. 2025. ”Oversight Structures for Agentic AI in Public-Sector Organizations.” https://arxiv.org/abs/2506.04836.
@@
Humble, Kristian. 2024. ”War, Artificial Intelligence, and the Future of Conflict.” https://gjia.georgetown.edu/2024/07/12/war-artificial-intelligence-and-the-future-of-conflict/.
@@
Pan, Alexander, Kush Bhatia, Jacob Steinhardt. 2022. ”The Effects of Reward Misspecification: Mapping and Mitigating Misaligned Models.” https://arxiv.org/abs/2201.03544.
@@
Hendrycks, Dan, Nicholas Carlini, John Schulman, Jacob Steinhardt. 2021. ”Unsolved Problems in ML Safety.” https://arxiv.org/pdf/2109.13916.
@@
Omohundro, Stephen M. 2008. ”The Basic AI Drives.” https://selfawaresystems.com/wp-content/uploads/2008/01/ai˙drives˙final.pdf.
@@
Wei, Jason, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, et al. 2022. ”Emergent Abilities of Large Language Models.” https://openreview.net/pdf?id=yzkSU5zdwD.
@@
Meinke, Alexander, Bronson Schoen, Jérémy Scheurer, Mikita Balesni, Rusheb Shah, et al. 2025. ”Frontier Models are Capable of In-context Scheming.” https://arxiv.org/abs/2412.04984.
@@
Rudner, Tim G. J., Helen Toner. 2021. ”Key Concepts in AI Safety: Specification in Machine Learning.” https://cset.georgetown.edu/wp-content/uploads/Key-Concepts-in-AI-Safety-Specification-in-Machine-Learning.pdf.
@@
Langosco di Langosco, Lauro, Jack Koch, Lee D. Sharkey, Jacob Pfau, David Krueger. 2022. ”Goal Misgeneralization in Deep Reinforcement Learning.” https://proceedings.mlr.press/v162/langosco22a/langosco22a.pdf.
@@
Fang, Richard, Rohan Bindu, Akul Gupta, Daniel Kang. 2024. ”LLM Agents can Autonomously Exploit One-day Vulnerabilities.” https://arxiv.org/pdf/2404.08144.
@@
Dalrymple, David, Joar Skalse, Yoshua Bengio, Stuart Russell, Max Tegmark, et al. 2024. ”Towards Guaranteed Safe AI: A Framework for Ensuring Robust and Reliable AI Systems.” https://arxiv.org/abs/2405.06624.
@@
Fang, Richard, Rohan Bindu, Akul Gupta, Qiusi Zhan, Daniel Kang. 2024. ”LLM Agents can Autonomously Hack Websites.” https://arxiv.org/abs/2402.06664.
@@
Christiano, Paul F., Jan Leike, Tom B. Brown, Miljan Martic, Shane Legg, et al. 2017. ”Deep Reinforcement Learning from Human Preferences.” https://dl.acm.org/doi/pdf/10.5555/3294996.3295184.
@@
Bai, Yuntao, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, et al. 2022. ”Training a Helpful and Harmless Assistant with Reinforcement Learning from Human Feedback.” https://arxiv.org/abs/2204.05862.
@@
Armstrong, Stuart, Xavier O’Rorke. 2018. ”Good and Safe Uses of AI Oracles.” https://arxiv.org/abs/1711.05541.
@@
Narayanan, Arvind, Sayash Kapoor. 2025. ”AI as Normal Technology.” https://knightcolumbia.org/content/ai-as-normal-technology.
@@
Grace, Katja, Harlan Stewart, Julia Fabienne Sandkühler, Stephen Thomas, Ben Weinstein-Raun, et al. 2025. ”Thousands of AI Authors on the Future of AI.” https://arxiv.org/pdf/2401.02843.
@@
Bai, Yuntao, Saurav Kadavath, Sandipan Kundu, Amanda Askell, Jackson Kernion, et al. 2022. ”Constitutional AI: Harmlessness from AI Feedback.” https://arxiv.org/abs/2212.08073.
@@
Turner, Alexander Matt, Logan Smith, Rohin Shah, Andrew Critch, Prasad Tadepalli. 2021. ”Optimal Policies Tend to Seek Power.” https://openreview.net/pdf?id=l7-DBWawSZH.
@@
Matheny, Jason. 2023. ”Artificial Intelligence: Challenges and Opportunities for the Department of Defense.” https://www.rand.org/content/dam/rand/pubs/testimonies/CTA2700/CTA2723-1/RAND˙CTA2723-1.pdf.
@@
Taylor, Jessica, Eliezer Yudkowsky, Patrick LaVictoire, Andrew Critch. 2016. ”Alignment for Advanced Machine Learning Systems.” https://intelligence.org/files/AlignmentMachineLearning.pdf.
@@
Askell, Amanda, Miles Brundage, Gillian Hadfield. 2019. ”The Role of Cooperation in Responsible AI Development.” https://arxiv.org/abs/1907.04534.
@@
Kasirzadeh, Atoosa. 2025. ”Two types of AI existential risk: decisive and accumulative.” https://link.springer.com/article/10.1007/s11098-025-02301-3.
@@
Clark, Jack, Dario Amodei. 2016. ”Faulty reward functions in the wild.” https://openai.com/index/faulty-reward-functions/.
@@
De Marzo, Giordano, Claudio Castellano, David Garcia. 2025. ”AI agents can coordinate beyond human scale.” https://arxiv.org/abs/2409.02822.
@@
Milli, Smitha, Dylan Hadfield-Menell, Anca Dragan, Stuart Russell. 2017. ”Should Robots be Obedient?” https://arxiv.org/abs/1705.09990.
@@
Armstrong, Stuart, Benjamin Levinstein. 2017. ”Low Impact Artificial Intelligences.” https://arxiv.org/abs/1705.10720.
@@
Boling, Bryan, Benjamin Boudreaux, Alexis A. Blanc, Christy Foran, Edward Geist. 2022. ”Emerging Technology Beyond 2035: Scenario-Based Technology Assessment for Future Military Contingencies.” https://www.rand.org/pubs/research˙reports/RRA1564-1.html.
@@
Bullock, Justin, Corin Katzke, Zershaaneh Qureshi, David Kristoffersson. 2024. ”AI Clarity: An Initial Research Agenda.” https://www.convergenceanalysis.org/publications/ai-clarity-an-initial-research-agenda.
@@
Qureshi, Zershaaneh. 2025. ”Pathways to short TAI timelines.” https://www.convergenceanalysis.org/research/pathways-to-short-tai-timelines.
@@
Katzke, Corin, Justin Bullock. 2024. ”AI governance needs a theory of victory.” https://www.convergenceanalysis.org/publications/ai-governance-needs-a-theory-of-victory.
@@
Katzke, Corin. 2024. ”Investigating the role of agency in AI x-risk.” https://www.convergenceanalysis.org/publications/investigating-the-role-of-agency-in-ai-x-risk.
@@
Katzke, Corin. 2024. ”Scenario planning for AI x-risk.” https://www.convergenceanalysis.org/publications/scenario-planning-for-ai-x-risk.
@@
Bullock, Justin, Elliot McKernon. 2024. ”Transformative AI and Scenario Planning for AI X-risk.” https://www.convergenceanalysis.org/publications/transformative-ai-and-scenario-planning-for-ai-x-risk.
@@
Dung, Leonard. 2024. ”Is superintelligence necessarily moral?” https://academic.oup.com/analysis/article-abstract/84/4/730/7774058?redirectedFrom=fulltext&login=false.
@@
Slattery, Peter, Alexander K. Saeri, Emily A. C. Grundy, Jess Graham, Michael Noetel. 2024. ”The AI Risk Repository: A Comprehensive Meta-Review, Database, and Taxonomy of Risks From Artificial Intelligence.” https://arxiv.org/abs/2408.12622.
@@
Hubinger, Evan, Chris van Merwijk, Vladimir Mikulik, Joar Skalse, Scott Garrabrant. 2021. ”Risks from Learned Optimization in Advanced Machine Learning Systems.” https://arxiv.org/abs/1906.01820v3.
@@
Gabriel, Iason, Arianna Manzini, Geoff Keeling, Lisa Anne Hendricks, Verena Rieser. 2024. ”The Ethics of Advanced AI Assistants.” https://storage.googleapis.com/deepmind-media/DeepMind.com/Blog/ethics-of-advanced-ai-assistants/the-ethics-of-advanced-ai-assistants-2024-i.pdf.
@@
Ji, Jiaming, Tianyi Qiu, Boyuan Chen, Borong Zhang, Hantao Lou. 2023. ”AI Alignment: A Comprehensive Survey.” https://arxiv.org/pdf/2310.19852.
AI Verify Foundation. 2023. ”Cataloguing LLM Evaluations.” https://aiverifyfoundation.sg/downloads/Cataloguing˙LLM˙Evaluations.pdf.
@@
Shevlane, Toby, Sebastian Farquhar, Ben Garfinkel, Mary Phuong, Jess Whittlestone. 2023. ”Model evaluation for extreme risks.” https://arxiv.org/pdf/2305.15324.
@@
Yudkowsky, Eliezer. 2008. ”Artificial Intelligence as a Positive and Negative Factor in Global Risk.” https://intelligence.org/files/AIPosNegFactor.pdf.
@@
Bradley, Patrick. 2020. ”Risk management standards and the active management of malicious intent in artificial superintelligence.” https://link.springer.com/article/10.1007/s00146-019-00890-2.
@@
Sotala, Kaj, Roman V. Yampolskiy. 2014. ”Responses to Catastrophic AGI Risk: A Survey.” https://iopscience.iop.org/article/10.1088/0031-8949/90/1/018001.
@@
Barrett, Anthony M., Seth D. Baum. 2017. ”A model of pathways to artificial superintelligence catastrophe for risk and decision analysis.” https://www.tandfonline.com/doi/full/10.1080/0952813X.2016.1186228.
@@
Amodei, Dario, Chris Olah, Jacob Steinhardt, Paul Christiano, John Schulman, et al. 2016. ”Concrete Problems in AI Safety.” https://arxiv.org/abs/1606.06565.
@@
Bommasani, Rishi, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, et al. 2022. ”On the Opportunities and Risks of Foundation Models.” https://arxiv.org/abs/2108.07258.
@@
Yampolskiy, Roman V. 2016. ”Taxonomy of Pathways to Dangerous Artificial Intelligence.” https://cdn.aaai.org/ocs/ws/ws0156/12566-57418-1-PB.pdf.
@@
Hogenhout, Lambert. 2021. ”A Framework for Ethical AI at the United Nations.” https://arxiv.org/pdf/2104.12547.
@@
Critch, Andrew, David Krueger. 2020. ”AI Research Considerations for Human Existential Safety (ARCHES).” https://arxiv.org/pdf/2006.04948.
@@
Sherman, Eli, Ian Eisenberg. 2024. ”AI Risk Profiles: A Standards Proposal for Pre-deployment AI Risk Disclosures.” https://ojs.aaai.org/index.php/AAAI/article/view/30348.
@@
McLean, Scott, Gemma J. M. Read, Jason Thompson, Chris Baber, Neville A. Stanton, et al. 2021. ”The Risks Associated with Artificial General Intelligence: A Systematic Review.” https://www.tandfonline.com/doi/full/10.1080/0952813X.2021.1964003#abstract.
@@
Steimers, André, Moritz Schneider. 2022. ”Sources of Risk of AI Systems.” https://www.mdpi.com/1660-4601/19/6/3641.
@@
Deng, Jiawen, Jiale Cheng, Hao Sun, Zhexin Zhang, Minlie Huang. 2023. ”Towards Safer Generative Language Models: A Survey on Safety Risks, Evaluations, and Improvements.” https://arxiv.org/pdf/2302.09270.
@@
Everitt, Tom, Gary Lea, Marcus Hutter. 2018. ”AGI Safety Literature Review.” https://arxiv.org/pdf/1805.01109.
@@
Tan, Samson, Araz Taeihagh, Kathy Baxter. 2022. ”The Risks of Machine Learning Systems.” https://arxiv.org/pdf/2204.09852.
@@
Hagendorff, Thilo. 2024. ”Mapping the Ethics of Generative AI: A Comprehensive Scoping Review.” https://arxiv.org/abs/2402.08323.
@@
Kilian, Kyle A., Christopher J. Ventura, Mark M. Bailey. 2022. ”Examining the Differential Risk from High-level Artificial Intelligence and the Question of Control.” https://arxiv.org/pdf/2211.03157.
@@
Weidinger, Laura, Jonathan Uesato, Maribeth Rauh, Conor Griffin, Po-Sen Huang. 2022. ”Taxonomy of Risks posed by Language Models.” https://dl.acm.org/doi/pdf/10.1145/3531146.3533088.
@@
Weidinger, Laura, Maribeth Rauh, Nahema Marchal, Arianna Manzini, Lisa Anne Hendricks. 2023. ”Sociotechnical Safety Evaluation of Generative AI Systems.” https://arxiv.org/pdf/2310.11986.
@@
Giarmoleo, Francesco Vincenzo, Ignacio Ferrero, Marta Rocchi, Massimiliano Matteo Pellegrini. 2024. ”What ethics can say on artificial intelligence: Insights from a systematic literature review.” https://onlinelibrary.wiley.com/doi/10.1111/basr.12336.
@@
Schnitzer, Ronald, Andreas Hapfelmeier, Sven Gaube, Hana Chockler, Hanna Wuerstle. 2024. ”AI Hazard Management: A framework for the systematic management of root causes for AI risks.” https://arxiv.org/pdf/2310.16727.
@@
Maas, Matthijs M. 2023. ”Advanced AI governance: A literature review of problems, options, and proposals.” https://law-ai.org/advanced-ai-gov-litrev/.
@@
Hadfield-Menell, Dylan, Anca Dragan, Pieter Abbeel, Stuart Russell. 2017. ”The Off-Switch Game.” https://arxiv.org/pdf/1611.08219.
@@
Founders Pledge. 2022. ”Autonomous weapon systems and military artificial intelligence (AI).” https://www.founderspledge.com/research/autonomous-weapon-systems-and-military-artificial-intelligence-ai.
@@
Denison, Carson, Monte MacDiarmid, Fazl Barez, David Duvenaud, Shauna Kravec. 2024. ”Sycophancy to Subterfuge: Investigating Reward-Tampering in Language Models.” https://arxiv.org/pdf/2406.10162.
@@
Krakovna, Victoria, Janos Kramar. 2023. ”Power-seeking can be probable and predictive for trained agents.” https://arxiv.org/pdf/2304.06528.
@@
Barnett, Peter, Lisa Thiergart. 2024. ”What AI evaluations for preventing catastrophic risks can and cannot do.” https://arxiv.org/html/2412.08653v1.
@@
Carlsmith, Joe. 2023. ”Scheming AIs: Will AIs fake alignment during training in order to get power?” https://arxiv.org/pdf/2311.08379.
@@
Uuk, Risto, Carlos Ignacio Gutierrez, Daniel Guppy, Lode Lauwaert, Atoosa Kasirzadeh. 2024. ”A Taxonomy of Systemic Risks from General-Purpose AI.” https://arxiv.org/pdf/2412.07780.
@@
Kinniment, Megan, Lucas Jun Koba Sato, Haoxing Du, Brian Goodrich, Max Hasin. 2024. ”Evaluating Language-Model Agents on Realistic Autonomous Tasks.” https://arxiv.org/pdf/2312.11671.

### v2

@@ 2.1 Literature Reviewed
Hastings-Woodhouse, Sarah. 2024. “Could we switch off a dangerous AI?” https://futureoflife.org/ai/could-we-switch-off-a-dangerous-ai/.
@@
Schmitz, Chris, Jonathan Rystrøm, Jan Batzner. 2025. “Oversight Structures for Agentic AI in Public-Sector Organizations.” https://arxiv.org/abs/2506.04836.
@@
Humble, Kristian. 2024. “War, Artificial Intelligence, and the Future of Conflict.” https://gjia.georgetown.edu/2024/07/12/war-artificial-intelligence-and-the-future-of-conflict/.
@@
Pan, Alexander, Kush Bhatia, Jacob Steinhardt. 2022. “The Effects of Reward Misspecification: Mapping and Mitigating Misaligned Models.” https://arxiv.org/abs/2201.03544.
@@
Hendrycks, Dan, Nicholas Carlini, John Schulman, Jacob Steinhardt. 2021. “Unsolved Problems in ML Safety.” https://arxiv.org/pdf/2109.13916.
@@
Omohundro, Stephen M. 2008. “The Basic AI Drives.” https://selfawaresystems.com/wp-content/uploads/2008/01/ai˙drives˙final.pdf.
@@
Wei, Jason, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, et al. 2022. “Emergent Abilities of Large Language Models.” https://openreview.net/pdf?id=yzkSU5zdwD.
@@
Meinke, Alexander, Bronson Schoen, Jérémy Scheurer, Mikita Balesni, Rusheb Shah, et al. 2025. “Frontier Models are Capable of In-context Scheming.” https://arxiv.org/abs/2412.04984.
@@
Rudner, Tim G. J., Helen Toner. 2021. “Key Concepts in AI Safety: Specification in Machine Learning.” https://cset.georgetown.edu/wp-content/uploads/Key-Concepts-in-AI-Safety-Specification-in-Machine-Learning.pdf.
@@
Langosco di Langosco, Lauro, Jack Koch, Lee D. Sharkey, Jacob Pfau, David Krueger. 2022. “Goal Misgeneralization in Deep Reinforcement Learning.” https://proceedings.mlr.press/v162/langosco22a/langosco22a.pdf.
@@
Fang, Richard, Rohan Bindu, Akul Gupta, Daniel Kang. 2024. “LLM Agents can Autonomously Exploit One-day Vulnerabilities.” https://arxiv.org/pdf/2404.08144.
@@
Dalrymple, David, Joar Skalse, Yoshua Bengio, Stuart Russell, Max Tegmark, et al. 2024. “Towards Guaranteed Safe AI: A Framework for Ensuring Robust and Reliable AI Systems.” https://arxiv.org/abs/2405.06624.
@@
Fang, Richard, Rohan Bindu, Akul Gupta, Qiusi Zhan, Daniel Kang. 2024. “LLM Agents can Autonomously Hack Websites.” https://arxiv.org/abs/2402.06664.
@@
Christiano, Paul F., Jan Leike, Tom B. Brown, Miljan Martic, Shane Legg, et al. 2017. “Deep Reinforcement Learning from Human Preferences.” https://dl.acm.org/doi/pdf/10.5555/3294996.3295184.
@@
Bai, Yuntao, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, et al. 2022. “Training a Helpful and Harmless Assistant with Reinforcement Learning from Human Feedback.” https://arxiv.org/abs/2204.05862.
@@
Armstrong, Stuart, Xavier O’Rorke. 2018. “Good and Safe Uses of AI Oracles.” https://arxiv.org/abs/1711.05541.
@@
Narayanan, Arvind, Sayash Kapoor. 2025. “AI as Normal Technology.” https://knightcolumbia.org/content/ai-as-normal-technology.
@@
Grace, Katja, Harlan Stewart, Julia Fabienne Sandkühler, Stephen Thomas, Ben Weinstein-Raun, et al. 2025. “Thousands of AI Authors on the Future of AI.” https://arxiv.org/pdf/2401.02843.
@@
Bai, Yuntao, Saurav Kadavath, Sandipan Kundu, Amanda Askell, Jackson Kernion, et al. 2022. “Constitutional AI: Harmlessness from AI Feedback.” https://arxiv.org/abs/2212.08073.
@@
Turner, Alexander Matt, Logan Smith, Rohin Shah, Andrew Critch, Prasad Tadepalli. 2021. “Optimal Policies Tend to Seek Power.” https://openreview.net/pdf?id=l7-DBWawSZH.
@@
Matheny, Jason. 2023. “Artificial Intelligence: Challenges and Opportunities for the Department of Defense.” https://www.rand.org/content/dam/rand/pubs/testimonies/CTA2700/CTA2723-1/RAND˙CTA2723-1.pdf.
@@
Taylor, Jessica, Eliezer Yudkowsky, Patrick LaVictoire, Andrew Critch. 2016. “Alignment for Advanced Machine Learning Systems.” https://intelligence.org/files/AlignmentMachineLearning.pdf.
@@
Askell, Amanda, Miles Brundage, Gillian Hadfield. 2019. “The Role of Cooperation in Responsible AI Development.” https://arxiv.org/abs/1907.04534.
@@
Kasirzadeh, Atoosa. 2025. “Two types of AI existential risk: decisive and accumulative.” https://link.springer.com/article/10.1007/s11098-025-02301-3.
@@
Clark, Jack, Dario Amodei. 2016. “Faulty reward functions in the wild.” https://openai.com/index/faulty-reward-functions/.
@@
De Marzo, Giordano, Claudio Castellano, David Garcia. 2025. “AI agents can coordinate beyond human scale.” https://arxiv.org/abs/2409.02822.
@@
Milli, Smitha, Dylan Hadfield-Menell, Anca Dragan, Stuart Russell. 2017. “Should Robots be Obedient?” https://arxiv.org/abs/1705.09990.
@@
Armstrong, Stuart, Benjamin Levinstein. 2017. “Low Impact Artificial Intelligences.” https://arxiv.org/abs/1705.10720.
@@
Boling, Bryan, Benjamin Boudreaux, Alexis A. Blanc, Christy Foran, Edward Geist. 2022. “Emerging Technology Beyond 2035: Scenario-Based Technology Assessment for Future Military Contingencies.” https://www.rand.org/pubs/research˙reports/RRA1564-1.html.
@@
Bullock, Justin, Corin Katzke, Zershaaneh Qureshi, David Kristoffersson. 2024. “AI Clarity: An Initial Research Agenda.” https://www.convergenceanalysis.org/publications/ai-clarity-an-initial-research-agenda.
@@
Qureshi, Zershaaneh. 2025. “Pathways to short TAI timelines.” https://www.convergenceanalysis.org/research/pathways-to-short-tai-timelines.
@@
Katzke, Corin, Justin Bullock. 2024. “AI governance needs a theory of victory.” https://www.convergenceanalysis.org/publications/ai-governance-needs-a-theory-of-victory.
@@
Katzke, Corin. 2024. “Investigating the role of agency in AI x-risk.” https://www.convergenceanalysis.org/publications/investigating-the-role-of-agency-in-ai-x-risk.
@@
Katzke, Corin. 2024. “Scenario planning for AI x-risk.” https://www.convergenceanalysis.org/publications/scenario-planning-for-ai-x-risk.
@@
Bullock, Justin, Elliot McKernon. 2024. “Transformative AI and Scenario Planning for AI X-risk.” https://www.convergenceanalysis.org/publications/transformative-ai-and-scenario-planning-for-ai-x-risk.
@@
Dung, Leonard. 2024. “Is superintelligence necessarily moral?” https://academic.oup.com/analysis/article-abstract/84/4/730/7774058?redirectedFrom=fulltext&login=false.
@@
Slattery, Peter, Alexander K. Saeri, Emily A. C. Grundy, Jess Graham, Michael Noetel. 2024. “The AI Risk Repository: A Comprehensive Meta-Review, Database, and Taxonomy of Risks From Artificial Intelligence.” https://arxiv.org/abs/2408.12622.
@@
Hubinger, Evan, Chris van Merwijk, Vladimir Mikulik, Joar Skalse, Scott Garrabrant. 2021. “Risks from Learned Optimization in Advanced Machine Learning Systems.” https://arxiv.org/abs/1906.01820v3.
@@
Gabriel, Iason, Arianna Manzini, Geoff Keeling, Lisa Anne Hendricks, Verena Rieser. 2024. “The Ethics of Advanced AI Assistants.” https://storage.googleapis.com/deepmind-media/DeepMind.com/Blog/ethics-of-advanced-ai-assistants/the-ethics-of-advanced-ai-assistants-2024-i.pdf.
@@
Ji, Jiaming, Tianyi Qiu, Boyuan Chen, Borong Zhang, Hantao Lou. 2023. “AI Alignment: A Comprehensive Survey.” https://arxiv.org/pdf/2310.19852.
AI Verify Foundation. 2023. “Cataloguing LLM Evaluations.” https://aiverifyfoundation.sg/downloads/Cataloguing˙LLM˙Evaluations.pdf.
@@
Shevlane, Toby, Sebastian Farquhar, Ben Garfinkel, Mary Phuong, Jess Whittlestone. 2023. “Model evaluation for extreme risks.” https://arxiv.org/pdf/2305.15324.
@@
Yudkowsky, Eliezer. 2008. “Artificial Intelligence as a Positive and Negative Factor in Global Risk.” https://intelligence.org/files/AIPosNegFactor.pdf.
@@
Bradley, Patrick. 2020. “Risk management standards and the active management of malicious intent in artificial superintelligence.” https://link.springer.com/article/10.1007/s00146-019-00890-2.
@@
Sotala, Kaj, Roman V. Yampolskiy. 2014. “Responses to Catastrophic AGI Risk: A Survey.” https://iopscience.iop.org/article/10.1088/0031-8949/90/1/018001.
@@
Barrett, Anthony M., Seth D. Baum. 2017. “A model of pathways to artificial superintelligence catastrophe for risk and decision analysis.” https://www.tandfonline.com/doi/full/10.1080/0952813X.2016.1186228.
@@
Amodei, Dario, Chris Olah, Jacob Steinhardt, Paul Christiano, John Schulman, et al. 2016. “Concrete Problems in AI Safety.” https://arxiv.org/abs/1606.06565.
@@
Bommasani, Rishi, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, et al. 2022. “On the Opportunities and Risks of Foundation Models.” https://arxiv.org/abs/2108.07258.
@@
Yampolskiy, Roman V. 2016. “Taxonomy of Pathways to Dangerous Artificial Intelligence.” https://cdn.aaai.org/ocs/ws/ws0156/12566-57418-1-PB.pdf.
@@
Hogenhout, Lambert. 2021. “A Framework for Ethical AI at the United Nations.” https://arxiv.org/pdf/2104.12547.
@@
Critch, Andrew, David Krueger. 2020. “AI Research Considerations for Human Existential Safety (ARCHES).” https://arxiv.org/pdf/2006.04948.
@@
Sherman, Eli, Ian Eisenberg. 2024. “AI Risk Profiles: A Standards Proposal for Pre-deployment AI Risk Disclosures.” https://ojs.aaai.org/index.php/AAAI/article/view/30348.
@@
McLean, Scott, Gemma J. M. Read, Jason Thompson, Chris Baber, Neville A. Stanton, et al. 2021. “The Risks Associated with Artificial General Intelligence: A Systematic Review.” https://www.tandfonline.com/doi/full/10.1080/0952813X.2021.1964003#abstract.
@@
Steimers, André, Moritz Schneider. 2022. “Sources of Risk of AI Systems.” https://www.mdpi.com/1660-4601/19/6/3641.
@@
Deng, Jiawen, Jiale Cheng, Hao Sun, Zhexin Zhang, Minlie Huang. 2023. “Towards Safer Generative Language Models: A Survey on Safety Risks, Evaluations, and Improvements.” https://arxiv.org/pdf/2302.09270.
@@
Everitt, Tom, Gary Lea, Marcus Hutter. 2018. “AGI Safety Literature Review.” https://arxiv.org/pdf/1805.01109.
@@
Tan, Samson, Araz Taeihagh, Kathy Baxter. 2022. “The Risks of Machine Learning Systems.” https://arxiv.org/pdf/2204.09852.
@@
Hagendorff, Thilo. 2024. “Mapping the Ethics of Generative AI: A Comprehensive Scoping Review.” https://arxiv.org/abs/2402.08323.
@@
Kilian, Kyle A., Christopher J. Ventura, Mark M. Bailey. 2022. “Examining the Differential Risk from High-level Artificial Intelligence and the Question of Control.” https://arxiv.org/pdf/2211.03157.
@@
Weidinger, Laura, Jonathan Uesato, Maribeth Rauh, Conor Griffin, Po-Sen Huang. 2022. “Taxonomy of Risks posed by Language Models.” https://dl.acm.org/doi/pdf/10.1145/3531146.3533088.
@@
Weidinger, Laura, Maribeth Rauh, Nahema Marchal, Arianna Manzini, Lisa Anne Hendricks. 2023. “Sociotechnical Safety Evaluation of Generative AI Systems.” https://arxiv.org/pdf/2310.11986.
@@
Giarmoleo, Francesco Vincenzo, Ignacio Ferrero, Marta Rocchi, Massimiliano Matteo Pellegrini. 2024. “What ethics can say on artificial intelligence: Insights from a systematic literature review.” https://onlinelibrary.wiley.com/doi/10.1111/basr.12336.
@@
Schnitzer, Ronald, Andreas Hapfelmeier, Sven Gaube, Hana Chockler, Hanna Wuerstle. 2024. “AI Hazard Management: A framework for the systematic management of root causes for AI risks.” https://arxiv.org/pdf/2310.16727.
@@
Maas, Matthijs M. 2023. “Advanced AI governance: A literature review of problems, options, and proposals.” https://law-ai.org/advanced-ai-gov-litrev/.
@@
Hadfield-Menell, Dylan, Anca Dragan, Pieter Abbeel, Stuart Russell. 2017. “The Off-Switch Game.” https://arxiv.org/pdf/1611.08219.
@@
Founders Pledge. 2022. “Autonomous weapon systems and military artificial intelligence (AI).” https://www.founderspledge.com/research/autonomous-weapon-systems-and-military-artificial-intelligence-ai.
@@
Denison, Carson, Monte MacDiarmid, Fazl Barez, David Duvenaud, Shauna Kravec. 2024. “Sycophancy to Subterfuge: Investigating Reward-Tampering in Language Models.” https://arxiv.org/pdf/2406.10162.
@@
Krakovna, Victoria, Janos Kramar. 2023. “Power-seeking can be probable and predictive for trained agents.” https://arxiv.org/pdf/2304.06528.
@@
Barnett, Peter, Lisa Thiergart. 2024. “What AI evaluations for preventing catastrophic risks can and cannot do.” https://arxiv.org/html/2412.08653v1.
@@
Carlsmith, Joe. 2023. “Scheming AIs: Will AIs fake alignment during training in order to get power?” https://arxiv.org/pdf/2311.08379.
@@
Uuk, Risto, Carlos Ignacio Gutierrez, Daniel Guppy, Lode Lauwaert, Atoosa Kasirzadeh. 2024. “A Taxonomy of Systemic Risks from General-Purpose AI.” https://arxiv.org/pdf/2412.07780.
@@
Kinniment, Megan, Lucas Jun Koba Sato, Haoxing Du, Brian Goodrich, Max Hasin. 2024. “Evaluating Language-Model Agents on Realistic Autonomous Tasks.” https://arxiv.org/pdf/2312.11671.